OpenStudy (anonymous):

What are the approximate solutions of 4x^2 + 3 = −12x to the nearest hundredth? Someone please show steps!!!!!!!1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (tkhunny):

Have you considered the Quadratic Formula?

OpenStudy (anonymous):

Yes I just ended up with whole numbers and my choice are all irrational decimals. I just need to see someone work it out

OpenStudy (tkhunny):

Did you use b = 12 (correct) or b = -12 (incorrect)?

OpenStudy (anonymous):

I used positive 12

OpenStudy (tkhunny):

Okay, then one piece at a time: 12^{2} - 4*4*3 = 144-48 = 96 \(\sqrt{96} = 4\sqrt{6}\) How about that? Pretty clear that you shouldn't have Whole Numbers after that.

OpenStudy (anonymous):

Where did you get the 4 and 6 and i dont inderstand how you got sqaure rrot of 96. I got the square root of 96/ 8 which is twelve?

OpenStudy (tkhunny):

I see: \(x = \dfrac{-12\pm\sqrt{96}}{8}\) is correct. This is the same as \(x = -\dfrac{12}{8}\pm\dfrac{\sqrt{96}}{8}\) This is NOT the same as \(x = -\dfrac{12}{8}\pm\sqrt{\dfrac{96}{8}}\) <== This is an error.

OpenStudy (anonymous):

I dont understand i=how to solve fromt here @tkhunny

OpenStudy (anonymous):

x ≈ −2.72 and x ≈ −0.28 I got it! this is rounded of course. Thanks!